The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #76

April 02, 1999 - 1:00 p.m. EST (10:00 a.m. PST)


The Lunar Prospector spacecraft continues to collect good science data in the extended mission orbit. Four of five science instruments are on: the Gamma Ray Spectrometer (GRS), the Neutron Spectrometer (NS), the Magnetometer (MAG) and the Electron Reflectometer (ER). Data continues to be collected for the Doppler Gravity Experiment (DGE) as well.

Investigation into the commanding anomalies during last week's maneuver sequence continues. There have been two commanding sequences since that maneuver for an attitude reorientation trim (DOY 088) and ER configuration commanding (DOY 089). The attitude sequence (17 cmds) was transmitted over DSS 61 with no anomalies. The ER sequence (19 cmds) was transmitted over DSS 16 and one command did not execute and had to be retransmitted. Both commanding sequences with failures were over 26-m DSN antenna sites, where the downlink telemetry also shows periods of noise.

The attitude reorientation timeline is given below (times are GMT, DOY 088):

1730 Clear maneuver registers

1734 Heaters ON 1740 GRS HV to 218 (not related to maneuver)

1742 Load maneuver parameters

1800 Fire 11 A1/A4 0.2 sec pulses (SEA change from -0.37 to 1.41)

1803 Safing

The attitude was precessed 1.4 deg; 0.03 kg of propellant was used.

Cycling of the transmitter during propellant tank heater events was discontinued on DOY 083. There have been 4 events where the battery voltage reached the yellow low voltage limit at the end of eclipse. Maximum eclipse occurs this weekend (DOY 093) and the decision was made to not reinstitute transmitter cycling but to closely monitor the battery voltage.

The Earth Moon Sensor and the Pressure Transducer were powered off at the end of the ER sequence (089/0446) to reduce the power load. Since then, there have been no battery yellow limit events. The ER commanding was to optimize data collection during the period of full moon, when the earth blocks the magnetic field of the sun and allows higher resolution of the lunar field.



Current spacecraft state (0000 GMT 4/2/99, DOY 092):

Orbit: 5501
Downlink: 3600 bps
Spin Rate: 12.09 rpm

Spin Axis Attitude (ecliptic):
Latitude: -88.6 deg
Longitude: 016 deg

Trajectory:
Periapsis Alt: 27.4 km
Apoapsis Alt: 32.6 km
Period: 111 min

Occultations: 49 minutes
Eclipses: 49 minutes
Propellant Remaining: 13.61 kg